The length of a rectangle is 6 yd longer than its width.
If the perimeter of the rectangle is 36 yd , find its area.


Answer by CharlesG2(834)

The length of a rectangle is 6 yd longer than its width.
If the perimeter of the rectangle is 36 yd , find its area
P perimeter = 2 * (L length + W width)
L=W+6
36=2*(W+6+W)
36=2*(2W+6)
18=2W+6
9=W+3
6=W
L=W+6
L=6+6
L=12
A=L*W=12*6=72 yards^2